OpenStudy (emmamink):

Solve for x: –4(3x – 2) = 6x + 2

OpenStudy (emmamink):

Can someone please help?

OpenStudy (anonymous):

-18x=-6 x=3

OpenStudy (anonymous):

wait

OpenStudy (cloverracer):

\[x = \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

-12x+8=6x+2 -18x= -6 x=-3

OpenStudy (emmamink):

Ok....

OpenStudy (emmamink):

Thanks! @CloverRacer

OpenStudy (emmamink):

It is correct CloverRacer! I appreciate your help!